A total of $58,000 is invested in two municipal bonds that pay 4.25% and 7.25% simple interest. The investor wants an annual interest income of $3335 from the investments. What amount should be invested in the 4.25% bond?
Answer by mananth(16946)

Part I 4.25% per annum ------------- Amount invested =x
Part II 7.25% per annum ------------ Amount invested = y
58000
Interest----- 3335.00

Part I 4.25% per annum ---x
Part II 7.25% per annum ---y
Total investment
x + 1 y= 58000 -------------1
Interest on both investments
4.25% x + 7.25% y= 3335
Multiply by 100
4.25 x + 7.25 y= 333500.00 --------2
Multiply (1) by -4.25
we get
-4.25 x -4.25 y= -246500.00
Add this to (2)
0 x 3 y= 87000
divide by 3
y = 29000
Part I 4.25% $ 29000
Part II 7.25% $ 29000

CHECK
29000 --------- 4.25% ------- 1232.50
29000 ------------- 7.25% ------- 2102.50
Total -------------------- 3335.00

m.ananth@hotmail.ca